the electron beam in a television tube consist of electron accelerated from rest through a potential difference of about 20000 v. what is the speed of the electrons?( ignore relativistic effects.) electron rest mass is 9.11x 10^31kg and electronic charge is 1.6x10^19c


1
Expert's answer
2022-03-14T13:06:11-0400

"E=\frac{mv^2}2,"

"\varphi=\frac Ee,"

"\varphi=\frac{mv^2}{2e},"

"v=\sqrt{\frac{2e\varphi}{m}},"

"v=84\cdot 10^6~\frac ms."